OpenStudy (anonymous):

What is the 100th term of the arithmetic sequence? 3, 8, 13, 18, … A. 300 B. 498 C. 500 D. 497

OpenStudy (igreen):

It's adding 5..

OpenStudy (igreen):

You can use this calculator to find the 100th term: http://www.basic-mathematics.com/arithmetic-sequence-calculator.html

OpenStudy (igreen):

100th term is 498!

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Natedogg1

OpenStudy (anonymous):

I got 300 tho lol?

OpenStudy (anonymous):

@igreen

OpenStudy (igreen):

No..it's 498..

OpenStudy (anonymous):

what did I do wrong to get 300?

OpenStudy (igreen):

http://www.basic-mathematics.com/arithmetic-sequence-calculator.html Do it yourself. In the first box, put in 3. Second box, 5. Third box, 100. 3 is the first term. 5 is the common difference. 100 is the term you want to find.

OpenStudy (igreen):

I don't know..what did you do to try to get the answer? @Natedogg1

OpenStudy (anonymous):

i got the common difference is 3 and sub 100-1 (n-1) and 3+99*3= 297+3 = 300

OpenStudy (igreen):

You thought the common difference was 3..but it's 5.

OpenStudy (igreen):

So it's: 3+99*5= 495+3 = 498

OpenStudy (anonymous):

the common is 5 i thought it was 3 ?

OpenStudy (anonymous):

ima try 300 answer see what happens lol

OpenStudy (anonymous):

lol your rite IGreen it was 498 xD

OpenStudy (igreen):

NO! DON'T! It's 5.. Look: 3, 8, 13, 18, … It's adding 5 everytime.. 3 + 5 = 8 Next term = 8 8 + 5 = 13 Next term = 13 13 + 5 = 18 Next term = 18 Etc..

OpenStudy (igreen):

c:

OpenStudy (anonymous):

so the second number in the pattern is the CD?= Common D lol